ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 34, 49, 93 Leon A. Petrosjan, St. Petersburg State University, Russia Time-consistancy problem in differential games N-person differential games are considered. There is a large variety of optmality principles (OP) definishing the behaviour of players in game. Many of them are implemented from classical (static) game theory. If some OP is chosen the game develops along corresponding optimal trajectory. It is impotant that the chosen OP remains optimal in subgames along the optimal trajectory (time consistency) Unfortunately this is not always the case. Regularization procedures are purposed to attain time consistent solutions.